# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 3.2 Evaluation of Informativeness: Representativeness
While uncertainty methods play a crucial role in deep AL, they still face certain challenges: 1. Outlier selection: The goal of using uncertainty in AL is to improve performance by querying hard samples of the current model. However, these methods could also select outliers that harm the model training (Karamcheti et al., 2021). This happens because uncertainty relies sorely on model predictions and ignores the exploration of intrinsic characteristics of the data distribution. Incorporating an additional informativeness measure to remove outliers would be beneficial for AL. 2. Distribution misalignment: Samples selected by uncertainty methods are often located near the decision boundary in the feature space (Settles, 2009). Therefore, the distribution of selected samples by uncertainty-based method may differ from the overall data distribution. This discrepancy may introduce dataset bias and lead to a performance drop. Therefore, the challenges above compel us further to discover the data distribution during the evaluation of informativeness.
Representativeness-based deep AL aims to select a subset of samples that can represent the entire dataset. Generally, highly representative samples are located in dense regions of the data manifold and contain information about other nearby samples. Moreover, these methods require diversity in sampling result. Representative samples should be widely distributed across the data manifold rather than concentrated in a specific region. Besides, representative samples should be visually distinctive in properties like imaging style or visual content. Deep feature representations encode such information and are used to calculate the mutual relationships between different samples. This section introduces three formulations of representativeness-based AL: cover-based, discrepancy-based, and density-based representativeness AL. The taxonomy of these methods is shown in Fig. 4.
@@ 3.2.1 Cover-based Active Learning
We can formulate representativeness-based AL as a problem of covering. A classic example of the covering problem is the facility location, such as covering all the city’s streets with some billboards (Farahani and Hekmatfar, 2009). Likewise, cover-based AL uses a few samples to cover the entire dataset. Ideally, these samples should be representative and contain information from other samples. These methods usually involve two settings: set cover and maximum coverage. Both settings are NP-hard, meaning they cannot be optimally solved in polynomial time. However, near-optimal solutions could be achieved in linear time using greedy algorithms. Specifically, the greedy algorithms iteratively select samples that cover other samples the most for annotation (Feige, 1998).
Set cover aims to select as few samples as possible to cover the entire dataset. CoreSet (Sener and Savarese, 2018) followed the setting of k-Center location (Hochbaum and Shmoys, 1985), which is also a variant of the set cover problem. In CoreSet, the L2 distance of deep features measures the similarity between different samples. They employed farthest-first traversal to solve the k-Center problem for selecting representative samples. Agarwal et al. (2020) introduced contextual diversity for AL, a metric that fused uncertainty and diversity of samples spatially and semantically. They replaced the L2 distance with contextual diversity and used the same method in CoreSet for sample selection. Caramalau et al. (2021) adopted graph convolutional networks (GCN) to model the relationships between labeled and unlabeled samples. GCNs improved the feature representation of unlabeled samples with the labeled dataset. Enhanced feature representation was further used for CoreSet sampling.
Maximum coverage selects a given number of samples to cover the entire dataset as much as possible. Yehuda et al. (2022) found that CoreSet tends to select outliers, especially when the annotation budget is low. To address this issue, they proposed ProbCover, which changed the setting from set cover to maximum coverage. They employed a graph-based greedy algorithm for sample selection. With the help of self-supervised deep features, ProbCover effectively avoided selecting outlier samples. Additionally, SA (Yang et al., 2017) provides another formulation of maximum coverage. SA first selected highly uncertain samples and then further chose representative samples for annotation. The representativeness was based on the cosine similarity of deep features. Specifically, sample $x$ is represented by the most similar sample from queried dataset $D_{t}^{q}$
@@
where a larger $R\left(D_{t}^{q},D_{t}^{u}\right)$ indicates that $D_{t}^{q}$ better represents $D_{t}^{u}$. It should be noted that SA is a generalization of the maximum coverage problem since the cosine similarity ranges from 0 to 1. But they still employed a greedy algorithm to find sample $x$ that maximizing $R\left(D_{t}^{q}\cup x,D_{t}^{u}\right)-R\left(D_{t}^{q},D_{t}^{u}\right)$. SA has inspired many subsequent works. Xu et al. (2018) quantized the segmentation networks in SA and found that it improved the accuracy of gland segmentation while significantly reducing memory usage. Zheng et al. (2019) proposed representative annotation (RA), which omits the uncertainty query in SA. RA trained a VAE for feature extraction and partitioned the feature space using hierarchical clustering. They selected representative samples in each cluster using a similar strategy to SA. Shen et al. (2020) changed the similarity measure in SA from $sim(\cdot,\cdot)$ to $1-sim(\cdot,\cdot)$, which enhanced the diversity of the selected samples. In keypoint detection of medical images, Quan et al. (2022) proposed a representative method to select template images for few-shot learning. First, they trained a feature extractor using self-supervised learning and applied SIFT for initial keypoint detection. Next, they calculated the average cosine similarity between template images and the entire dataset. Finally, they picked the template combination with the highest similarity for annotation.
@@
View of submodular functions: Both set cover and maximum coverage can be formulated from the perspective of submodular functions (Fujishige, 2005). These functions show diminishing marginal returns, meaning each added element brings less gain than the previous one as the set gets larger. Generally, each submodular function corresponds to a particular optimization problem. If a submodular function is monotonic and non-negative, we can use a greedy algorithm to get near-optimal solutions in linear time. In cover-based AL, methods like SA and RA followed the setting of submodular functions, but the authors didn’t present their methods from this perspective. Introducing submodular functions would extend the formulation of AL and ensure the selected samples are both representative and diverse. Typical steps for this type of method involve calculating sample similarities, constructing a submodular optimization problem, and solving it using a greedy algorithm (Wei et al., 2015). Kothawade et al. (2021) introduced an AL framework based on submodular information measures, effectively addressing issues such as scarcity of rare class, redundancy, and out-of-distribution data. In object detection, Kothawade et al. (2022a) focused on samples of minority classes. They constructed a set of classes of interest and selected unlabeled samples similar to these classes for annotation through submodular mutual information (Kothawade et al., 2022b).
@@ 3.2.2 Discrepancy-based Active Learning
In discrepancy-based AL, unlabeled samples farthest from the labeled set are considered the most representative. The main idea is that if we queried such samples for multiple rounds, the discrepancy between the distributions of labeled and unlabeled sets would be significantly reduced. Therefore, a small set of samples could well represent the entire dataset. The key to these methods is measuring the discrepancy (i.e., distance) between two high-dimensional distributions. This section presents three metrics for measuring discrepancy: similarity-based discrepancy, H-divergence, and Wasserstein distance.
@@
Similarity-based discrepancy: As a practical and easy-to-implement metric, we can approximate the distance between distributions based on sample similarity. Caramalau et al. (2021) proposed UncertainGCN, which employed GCN to model the relationship between labeled and unlabeled samples. They selected the unlabeled samples with the lowest similarity to the labeled set. In gland and MRI infant brain segmentation, Li and Yin (2020) adopted the average cosine similarity as the distance between two datasets. They selected samples far from the labeled set and close to the unlabeled set. In object detection, Wu et al. (2022a) constructed prototypes with sample features and prediction entropy. They selected unlabeled samples that were far from the labeled prototype.
@@
H-divergence estimates the distance of distribution with the help of the discriminator from generative adversarial networks (GAN) (Goodfellow et al., 2014a). More specifically, the discriminator tries to distinguish between labeled and unlabeled samples, and there is a close relationship between H-divergence and the discriminator’s output (Gissin and Shalev-Shwartz, 2019). Variational adversarial active learning (VAAL) (Sinha et al., 2019) combined VAE with a discriminator for discrepancy-based AL. In VAAL, the VAE mapped samples to a latent space while the discriminator distinguished whether samples were labeled. These two are mutually influenced by adversarial training. VAE tried to fool the discriminator into judging all samples as labeled while the discriminator attempted to correctly differentiate between labeled and unlabeled samples. After multiple rounds of adversarial training, VAAL selected samples that the discriminator deemed most likely to be unlabeled for annotation. Unlike VAAL, Gissin and Shalev-Shwartz (2019) trained the discriminator without adversarial training. Zhang et al. (2020) replaced the discriminator’s binary label with sample uncertainty. They also combined features of VAE with features from the supervised model. Wang et al. (2020b) adopted a neural network module for sample selection. To train such a module, they added another discriminator on top of VAAL, which aimed to differentiate between the real and VAE-reconstructed features for unlabeled samples. After adversarial training of both discriminators, the module selected uncertain and representative samples. Kim et al. (2021) combined LL4AL with VAAL, feeding both loss ranking predictions and VAE features into the discriminator.
@@
Wasserstein distance is widely used for computing distribution distances. Shui et al. (2020) indicated that H-divergence may compromise the diversity of sample selection, while Wasserstein distance ensures the queried samples are representative and diverse. They further proposed Wasserstein adversarial active learning (WAAL). Specifically, WAAL was built upon VAAL and adopted an additional module for sample selection. They trained this module by minimizing the Wasserstein distance between labeled and unlabeled sets. WAAL selected samples that are highly uncertain and most likely to be unlabeled for annotation. Mahmood et al. (2022) formulated AL as an optimal transport problem. They aimed at minimizing the Wasserstein distance between the labeled and unlabeled sets with self-supervised features. They further adopted mixed-integer programming that guarantees global convergence for diverse sample selection. Moreover, Xie et al. (2023b) considered the candidates as continuously optimizable variables based on self-supervised features. They randomly initialized the candidate samples at first. Then, they maximized the similarity between candidates and their nearest neighbors while minimizing the similarity between candidates and labeled samples. Finally, they selected the nearest neighbors of the final candidates for annotation. They proved the objective is equivalent to minimizing the Wasserstein distance between the labeled and unlabeled samples.
@@ 3.2.3 Density-based Active Learning
Density-based AL employs density estimation to characterize the data distribution in a high-dimensional feature space. The likelihood is the estimated density of the data distribution, and a more densely populated area indicates a higher likelihood. In this case, representative samples are samples with high likelihood. However, such methods can easily cause redundancy in sample selection. As a result, techniques like clustering are frequently used to improve diversity in sample selection. Density-based AL directly estimates the data distribution, which prevents the need to solve complex optimization problems. TypiClust (Hacohen et al., 2022) projected samples to a high-dimensional feature space via a self-supervised encoder. The density of a sample was defined as the reciprocal of the L2 distances to its k-nearest neighbors. Additionally, TypiClust performed clustering beforehand to ensure the diversity of selected samples. Wang et al. (2022c) proposed two variants of density-based AL. The first variant fixed the feature representation. The process was similar to TypiClust, but they maximized the distances between selected samples to ensure diversity. The other variant was in an end-to-end fashion. Feature representation and sample selection were trained simultaneously. This variant used a learnable k-Means clustering to jointly optimize cluster assignment and feature representation with a local smoothness constraint.
In active domain adaptation, density estimation is also widely used to select representative samples in the target domain. Please refer to §4.3 for related works.

### v3

@@ 3.2 Evaluation of Informativeness: Representativeness
While uncertainty-based methods play a crucial role in deep AL, they still face certain challenges: 1. Outlier selection: The goal of using uncertainty in AL is to improve performance by querying hard samples of the current model. However, these methods could also select outliers that harm the model training (Karamcheti et al., 2021). This happens mainly because uncertainty-based methods often ignore the intrinsic characteristics of the sample itself. 2. Distribution misalignment: In the feature space, uncertain samples are often located near the decision boundary (Settles, 2009). Therefore, the distribution of samples selected by uncertainty-based methods is usually different from the overall data distribution. This discrepancy introduces dataset bias and leads to a performance drop. This challenge could be alleviated if the relationship between different samples is carefully considered during the AL query. In summary, uncertainty-based AL lacks exploration of the visual information carried in each sample and the relationship between different samples. These challenges above call for a new informativeness metric in AL.
Representativeness is adopted in AL to overcome challenges brought by uncertainty. Representativeness-based AL aims to select a subset of samples that can represent the entire dataset. Specifically, representative samples should be visually distinctive in properties like imaging style or visual content. In medical image analysis, images are often high-dimensional and thus computation-intensive for the DL model. Besides, important information like lesions or tissues is not always directly visible or easily distinguishable. A good feature representation greatly reduces the image dimensionality and also extracts anatomical, histological, pathological, or even functional information in medical images. Therefore, the query process of representativeness-based AL is often conducted in the feature space. Besides, representative samples should also be widely distributed across the data distribution rather than concentrated in a specific region. In other words, these samples should be diverse. This is to minimize redundancy in the query result and try to keep the original data distribution as much as possible. Therefore, proper metrics of sample-wise or distribution-wise distance and sample density are needed to assess the landscape of a dataset. Besides, the uniqueness of medical images may require different distance metrics than that of natural images. This section introduces four types of representativeness-based AL: clustering-based, cover-based, discrepancy-based, and density-based representativeness AL. The taxonomy of these methods is shown in Fig. 4.
@@
3.2.1 Clustering-based Active Learning
With the advancement of feature extraction in medical image analysis, images with similar appearances tend to group together in the feature space (Zheng et al., 2019). Therefore, a straightforward approach involves clustering on the data embeddings to select representative samples. This type of method grouped the data into several clusters and then selected the centroid samples of each cluster. It leveraged the inherent structure within the data for insightful groupings and was also very easy to implement. K-Means was the most popular choice in clustering-based active learning. Pourahmadi et al. (2021) performed k-Means on the off-the-shelf self-supervised features, then selected cluster centers for annotation. Based on the self-supervised feature, Jin et al. (2022b) adopted the k-Means++ for clustering and the silhouette coefficient to determine the optimal number of clusters. Their proposed method achieved commendable performance in lung segmentation of chest X-rays and lesion segmentation of dermoscopic images. In nuclei segmentation, Lou et al. (2023) performed coarse-level and fine-level clustering using K-Means, aiming to select informative patches from pathological images. In connectomics, Lin et al. (2020) proposed two-stream clustering for active selection. They first predicted the semantic mask for each unlabeled sample and simplified the AL task to judge the correctness of each predicted ROI. Besides, they trained two feature extractors of VAE with segmentation masks and unlabeled images, respectively. For two-stream clustering, they first applied mask-level clustering with mask features to group ROIs with similar appearances. Within each mask cluster, image-level clustering is further performed. This method achieved excellent performance in synapse detection and mitochondria segmentation. Results also showed that two-stream clustering outperforms clustering with concatenated mask and image features by preventing the image feature from dominating the results.
@@ 3.2.2 Cover-based Active Learning
We can formulate representativeness-based AL as a covering problem. A classic example of the covering problem is facility location, such as covering all the city’s streets with billboards (Farahani and Hekmatfar, 2009). Likewise, cover-based AL uses a few samples to cover the entire dataset, which is analogous to using several spheres to cover all the samples in the feature space, with the center of each sphere being the selected sample. Ideally, these samples should be representative and contain information on other samples. These methods usually involve two settings: set cover and maximum cover. Both settings are NP-hard, meaning they cannot be optimally solved in polynomial time. However, near-optimal solutions could be achieved in linear time using greedy algorithms, which iteratively select samples that cover most of the other samples for annotation (Feige, 1998). These two variants are slightly different in the problem setting. The set cover is constrained by complete coverage, which means that it cannot omit any sample in the dataset. To achieve this goal, the radius of its covering spheres may be very large, and when there are very few samples selected, outliers might be chosen as the centers of the spheres (Yehuda et al., 2022). The goal of max cover is to cover as much of the entire dataset as possible, breaking the constraint of set cover, and thus avoiding the issue of outlier selection.
Set cover: Core-Set (Sener and Savarese, 2018) followed the setting of k-Center location (Hochbaum and Shmoys, 1985), which is also a variant of the set cover problem. They employed farthest-first traversal to solve the k-Center problem for selecting representative samples. The L2 distance of deep features is used to measure the similarity between different samples. Agarwal et al. (2020) introduced contextual diversity for AL, a metric that fused uncertainty and diversity of samples spatially and semantically. They replaced the L2 distance with contextual diversity and adopted the same farthest-first traversal for sample selection as Sener and Savarese (2018). Caramalau et al. (2021) adopted graph convolutional networks (GCN) to model the relationships between labeled and unlabeled samples. GCNs improved the feature representation of unlabeled samples with the labeled dataset. Enhanced feature representation was further used for Core-Set sampling.
Maximum cover: As a pioneering work of AL in medical image analysis, SA (Yang et al., 2017) stands out as one of the initial endeavors to introduce the concept of representativeness into AL. SA first selected highly uncertain samples and then chose representative samples for annotation. The formulation of the representativeness part in SA followed the setting of maximum cover. The representativeness metric was based on the cosine similarity of deep features. Specifically, sample $x$ is represented by the most similar sample from queried dataset $D_{t}^{q}$
@@
where a larger $R\left(D_{t}^{q},D_{t}^{u}\right)$ indicates that $D_{t}^{q}$ better represents $D_{t}^{u}$. It should be noted that SA is a generalization of the maximum cover problem since the cosine similarity ranges from 0 to 1. But they still employed a greedy algorithm to find sample $x$ that maximizing $R\left(D_{t}^{q}\cup x,D_{t}^{u}\right)-R\left(D_{t}^{q},D_{t}^{u}\right)$. Many subsequent AL works built their framework of cover-based AL on SA, especially in the field of medical image analysis. Xu et al. (2018) quantized the segmentation networks in SA and found that it improved the accuracy of gland segmentation while significantly reducing memory usage. Zheng et al. (2019) proposed representative annotation (RA), which omits the uncertainty query in SA. RA trained a VAE for feature extraction and partitioned the feature space using hierarchical clustering. They selected representative samples in each cluster using a similar strategy to SA. RA achieved superior performance in gland segmentation on histological images, fungus segmentation on electron microscopy images, and whole heart segmentation on MRI. In breast cancer segmentation on immunohistochemistry images, Shen et al. (2020) changed the similarity measure in SA from $sim(\cdot,\cdot)$ to $1-sim(\cdot,\cdot)$, which enhanced the diversity of the selected samples. Additionally, some works follow different formulations than those of SA in maximum cover. In keypoint detection of medical images, Quan et al. (2022) proposed a representative method to select template images for few-shot learning. First, they trained a feature extractor using self-supervised learning and applied the scale-invariant feature transform descriptor for initial keypoint detection. Next, they calculated the average cosine similarity between template images and the entire dataset. Finally, they picked the template combination with the highest similarity for annotation. Yehuda et al. (2022) found that Core-Set (Sener and Savarese, 2018), which followed the setting of the set cover, tends to select outliers, especially when the annotation budget is low. To address this issue, they proposed ProbCover, which changed the setting from set cover to maximum cover. With the help of self-supervised deep features and a graph-based greedy algorithm, ProbCover effectively avoided outlier selection in cover-based AL.
@@
View of submodular functions: Both set cover and maximum cover can be formulated from the perspective of submodular set functions (Fujishige, 2005). These functions show diminishing returns. Specifically, given two sets $A$ and $B$, $A\subset B$, for every element $z$ that not in $B$, a submodular set function $g$ has that $g\left(A\cup{z}\right)-g\left(A\right)\geq g\left(B\cup{z}\right)-g\left(B\right)$. This property makes submodular set functions suitable for AL. Suppose the informativeness function $I$ is submodular. It means that each newly queried sample brings less informativeness gain than the previous one, which indicates that highly informative samples should be queried first. Besides, if we can formulate optimization problems in terms of monotonic and submodular functions, we can use a greedy algorithm to get near-optimal solutions in linear time. For AL, if $I$ is submodular and monotonic, it means that we could greedily select the samples that maximize $I$. In cover-based AL, methods like SA and RA followed the setting of submodular functions, but the authors didn’t present their methods from this perspective. Introducing submodular functions would extend the formulation of AL and ensure the selected samples are both representative and diverse. Typical steps for this type of method involve calculating sample similarities, constructing a submodular optimization problem, and solving it using a greedy algorithm (Wei et al., 2015). Kothawade et al. (2021) introduced an AL framework based on submodular information measures, effectively addressing issues such as scarcity of rare class, redundancy, and out-of-distribution data. In object detection, Kothawade et al. (2022a) focused on samples of minority classes. They first constructed a reference dataset containing samples of certain classes of interest. Then, unlabeled samples similar to the reference set for annotation through submodular mutual information (SMI). SMI is used to measure the similarity between two sets. Suppose two sets $A$, $B$ and a submodular function $g$, the SMI is defined as $\mathcal{I}_{SMI}=g(A)+g(B)-g(A\cup B)$. Please refer to Kothawade et al. (2022b) for more detailed definitions of SMI.
@@ 3.2.3 Discrepancy-based Active Learning
In discrepancy-based AL, unlabeled samples farthest from the labeled set are considered the most representative. The main idea is that if we queried such samples for multiple rounds, the discrepancy between the distributions of labeled and unlabeled sets would be significantly reduced. Therefore, a small set of samples could well represent the entire dataset. The key to these methods is measuring the discrepancy (i.e., distance) between two high-dimensional distributions. In this section, we present four discrepancies between probability distributions: similarity-based discrepancy, H-divergence, Wasserstein distance, and maximum mean discrepancy (MMD).
@@
Similarity-based discrepancy: As a practical and easy-to-implement metric, we can approximate the distance between distributions based on sample similarity. In gland and MRI infant brain segmentation, Li and Yin (2020) adopted the average cosine similarity as the distance between two datasets. They selected samples far from the labeled set and close to the unlabeled set. Caramalau et al. (2021) proposed UncertainGCN, which employed GCN to model the relationship between labeled and unlabeled samples. They selected the unlabeled samples with the lowest similarity to the labeled set. In object detection, Wu et al. (2022a) constructed prototypes with sample features and prediction entropy. They selected unlabeled samples that were far from the labeled prototype.
@@
H-divergence estimates the distance of distribution with the help of the discriminator from generative adversarial networks (GAN) (Goodfellow et al., 2014a). More specifically, the discriminator tries to distinguish between labeled and unlabeled samples, and there is a close relationship between H-divergence and the discriminator’s output (Gissin and Shalev-Shwartz, 2019). Variational adversarial active learning (VAAL) (Sinha et al., 2019) combined VAE with a discriminator for discrepancy-based AL. In VAAL, the VAE mapped samples to a latent space while the discriminator distinguished whether samples were labeled. These two are mutually influenced by adversarial training. VAE tried to fool the discriminator into judging all samples as labeled while the discriminator attempted to differentiate between labeled and unlabeled samples correctly. After multiple rounds of adversarial training, VAAL selected samples that the discriminator deemed most likely to be unlabeled for annotation. VAAL inspired many subsequent works. Khanal et al. (2023) adopted multimodal information to improve VAAL. For multimodal medical images, they modified the VAE to reconstruct images of both modalities using the latent code of only one modality. The proposed method was evaluated on brain tumor segmentation, classification, and chest X-ray classification. Gissin and Shalev-Shwartz (2019) trained the discriminator without adversarial training. Zhang et al. (2020) replaced the discriminator’s binary label with sample uncertainty. They also combined features of VAE with features from the supervised model. Wang et al. (2020c) adopted a neural network module for sample selection. To train such a module, they added another discriminator on top of VAAL, which aimed to differentiate between the real and VAE-reconstructed features for unlabeled samples. After adversarial training of both discriminators, the module selected uncertain and representative samples. Kim et al. (2021) combined learning loss for active learning with VAAL, feeding both loss ranking predictions and VAE features into the discriminator.
@@
Wasserstein distance is widely used for computing distribution distances. Shui et al. (2020) indicated that H-divergence compromises the diversity of sample selection, while Wasserstein distance ensures the queried samples are representative and diverse. They further proposed Wasserstein adversarial active learning (WAAL), which built upon VAAL and adopted an additional module for sample selection. They trained this module by minimizing the Wasserstein distance between labeled and unlabeled sets. WAAL selected samples that are highly uncertain and most likely to be unlabeled for annotation. Mahmood et al. (2022) formulated AL as an optimal transport problem. They aimed at minimizing the Wasserstein distance between the labeled and unlabeled sets with self-supervised features. They further adopted mixed-integer programming that guarantees global convergence for diverse sample selection. Moreover, Xie et al. (2023b) considered the candidates as continuously optimizable variables based on self-supervised features. They randomly initialized the candidate samples at first. Then, they maximized the similarity between candidates and their nearest neighbors while minimizing the similarity between candidates and labeled samples. Finally, they selected the nearest neighbors of the final candidates for annotation. They proved the objective is equivalent to minimizing the Wasserstein distance between the labeled and unlabeled samples.
Maximum mean discrepancy measures the distance of two distributions as the distances between their mean features with kernel trick (Gretton et al., 2012). In active domain adaptation (will be detailed in §4.5), Hwang et al. (2022) adopted MMD to measure the distance between the source and target domain. Then, MMD was used to select representative and diverse samples in the target domain. It should be noted that the Wasserstein distance belongs to the family of integral probability metrics (IPM), while MMD simultaneously falls into the range of IPM and previously mentioned H-divergence. Please refer to Zhao et al. (2022) for a more detailed taxonomy of the discrepancy between probability distributions.
@@ 3.2.4 Density-based Active Learning
Density-based active learning tends to select samples from the most densely populated area of the data distribution. It employs density estimation to characterize the data distribution in a high-dimensional feature space. The likelihood is the estimated density of the data distribution, and a more densely populated area indicates a higher likelihood. In this case, representative samples are samples with high likelihood. However, such methods can easily cause redundancy in sample selection. As a result, techniques like clustering are frequently used to improve diversity in sample selection. Density-based AL directly estimates the data distribution, which prevents the need to solve complex optimization problems. In shoulder MRI musculoskeletal segmentation, Ozdemir et al. (2021) adopted infoVAE (Zhao et al., 2017) to estimate the density of each sample in the labeled dataset and unlabeled pool. Specifically, MMD replaced the KL divergence as the regularization term in the training of infoVAE. The posterior probability by the encoder was used as the density metric. Samples with higher density regarding the unlabeled pool and lower density regarding the labeled dataset were selected for annotation. TypiClust (Hacohen et al., 2022) projected samples to a high-dimensional feature space via a self-supervised encoder. The density of a sample was defined as the reciprocal of the L2 distances to its k-nearest neighbors. Additionally, TypiClust performed clustering beforehand to ensure the diversity of selected samples. Wang et al. (2022c) proposed two variants of density-based AL. The first variant fixed the feature representation. The process was similar to TypiClust, but they maximized the distances between selected samples to ensure diversity. The other variant was in an end-to-end fashion. Feature representation and sample selection were trained simultaneously. This variant used a learnable k-Means clustering to jointly optimize cluster assignment and feature representation with a local smoothness constraint.
It is worth noting that cover-based and density-based AL differ in both concept and methodology. In concept, samples in cover-based AL tend to cover the entire dataset. However, they do not have to lie in the densest area of the data distribution. For example, Yehuda et al. (2022) showed that Core-Set (Sener and Savarese, 2018), a popular cover-based method, tends to select outliers in the low-budget regime. In this case, cover-based AL is opposite to density-based AL, which also indicates that density-based AL may be a better choice in the low-budget regime. From the perspective of methodology, cover-based AL needs to solve an NP-hard problem in linear time with a greedy algorithm. Although this algorithm results in acceptable solutions, it’s almost impossible to know how the AL performance would be if the optimal solutions could be achieved. For density-based AL, the NP-hard problem is replaced with density estimation, which is more computation-efficient.